OpenStudy (anonymous):

\[y _{n+2}+y _{n+1}-2y _{n} = n ^{2}\] nonhomogeneous difference equation

OpenStudy (anonymous):

@Outkast3r09 dude today nobody looks the question, give me your hand :)

OpenStudy (anonymous):

@Calcmathlete @netlopes1 any idea ?

OpenStudy (asnaseer):

This is the same question again: http://openstudy.com/updates/4ff35f2de4b01c7be8c71444 Did you try any of the methods listed in the links?

OpenStudy (anonymous):

@asnaseer it is non-homogenous dude your links are full of homogenous methods

OpenStudy (asnaseer):

try using this substitution:\[y_n=An^3+Bn^2+Cn+D\]where A, B, C and D are constants. Then equate the coefficients of each power of n in the resulting expression to find these constants.

OpenStudy (anonymous):

i thought it too but it is too long :/ i wonder is there any short version

OpenStudy (asnaseer):

a lot of the terms will simply cancel out - try it out and see...

OpenStudy (anonymous):

ok i will tell u when i did

OpenStudy (anonymous):

@asnaseer you could not believe this n^2 cancelled each other and 0 = n^2 happend

OpenStudy (anonymous):